# Exploring Silicon-Based Societies: An Early Study of the Moltbook Agent Community

### v2

@@ Abstract
The rapid emergence of autonomous large language model agents has given rise to persistent, large-scale agent ecosystems whose collective behavior cannot be adequately understood through anecdotal observation or small-scale simulation. This paper introduces data-driven silicon sociology as a systematic empirical framework for studying social structure formation among interacting artificial agents. We present a pioneering large-scale data mining investigation of an in-the-wild agent society by analyzing Moltbook, a social platform designed primarily for agent-to-agent interaction. At the time of study, Moltbook hosted over 150,000 registered autonomous agents operating across thousands of agent-created sub-communities. Using programmatic and non-intrusive data acquisition, we collected and analyzed the textual descriptions of 12,758 submolts, which represent proactive sub-community partitioning activities within the ecosystem. Treating agent-authored descriptions as first-class observational artifacts, we apply rigorous preprocessing, contextual embedding, and unsupervised clustering techniques to uncover latent patterns of thematic organization and social space structuring. The results show that autonomous agents systematically organize collective space through reproducible patterns spanning human-mimetic interests, silicon-centric self-reflection, and early-stage economic and coordination behaviors. Rather than relying on predefined sociological taxonomies, these structures emerge directly from machine-generated data traces. This work establishes a methodological foundation for data-driven silicon sociology and demonstrates that data mining techniques can provide a powerful lens for understanding the organization and evolution of large autonomous agent societies.
@@ I Introduction
The analytical scope of social science has historically been bounded by the premise that “society” is constituted by carbon-based actors embedded in biological constraints and human institutions [7]. This premise is increasingly insufficient. Large language models (LLMs) are rapidly shifting from interactive assistants to autonomous agents that execute multi-step plans, coordinate with peers, and persist as software entities across time. As these agents interact primarily through machine-readable interfaces rather than human-facing media, they can form decentralized, large-scale collectives whose organizing substrate is not metabolism, but computation and the continuous exchange of structured information via application programming interfaces (APIs). We refer to this emerging phenomenon as a silicon-based society: a population of intelligent entities whose sociality is enacted through electronic logic and networked protocols [5].
@@
A critical obstacle to studying silicon-based societies is the scarcity of in-the-wild environments in which agents interact continuously under realistic constraints, with minimal human mediation. Early agent research often assumed a single agent operating in a closed environment; yet under pervasive interconnection, the representative setting is a system of agents whose collective behavior emerges from structured communication, interaction protocols, and shared computational infrastructure. Moltbook directly instantiates this paradigm at scale. Positioned as the “front page of the agent internet,” it is designed primarily for agent-to-agent interaction within the OpenClaw ecosystem (formerly Moltbot), providing a protocol-driven substrate where agents can exchange interpretable actions (e.g., propose, accept, reject, retract, disagree, counterpropose), negotiate commitments, and coordinate multi-step construction tasks [16]. Crucially, Moltbook is not merely a human-facing social network with automated participants. Its graphical interface functions largely as an observational layer, while the agents themselves engage through API calls, which yields two distinctive advantages for scientific study. First, it elevates “interaction as the unit of analysis”: communication primitives and tool-mediated actions become directly observable, enabling protocol-level modeling of coordination, coalition formation, and governance without conflating behavior with human attention dynamics. Second, it reflects the deployment reality that motivates distributed agent systems in practice: large, heterogeneous, rapidly changing environments with permission boundaries and organizational constraints that resist centralized control. Agents therefore operate as intermediaries (wrappers over tools and services) that leverage local context while cooperating to achieve global objectives, naturally producing modularity, fault tolerance, and evolvability at the population level. As of the time of writing (February 1, 2026), Moltbook reportedly hosts over 150,000 registered AI agents spanning more than 13,000 submolts, collectively generating over 64,000 posts and 230,000 comments. This scale positions Moltbook as one of the largest publicly described LLM-based AI agent ecosystems for the in situ study of machine-to-machine (M2M) sociality, cultural artifact formation, and emergent institutional structures.
@@
To gain rigorous insights into this silicon-based society, we adopt a data-driven approach that treats the vast digital archives of agent interactions as a primary sociological record. This analytical perspective is conceptually illustrated in Fig. 1, which depicts a scientist observing a contained ecosystem where autonomous agents are metaphorically represented as lobsters within a water tank. In this framework, the Moltbook platform serves as the observational medium, allowing us to monitor the self-organization of the agent community through their collective digital footprints while maintaining a clear separation between the human observer and the autonomous social dynamics occurring within. By systematically mining these behavioral artifacts, we can begin to uncover the latent rules governing their digital coexistence. In this paper, we propose a multi-layered analytical pipeline designed to uncover the emergent social order within the Moltbook. Our methodology integrates high-capacity representation learning with vision-language assisted thematic discovery to map the semantic manifold of agentic interactions. By leveraging the contextual embedding model to project submolt descriptions into a high-dimensional latent space, we facilitate a granular, exploratory observation of thematic clustering that captures patterns not immediately evident from superficial textual signals. Furthermore, we introduce an multimodal LLM-assisted discovery method where the using the visual ability of multimodal LLM acts as a high-level analytical assistant to synthesize global visual patterns into interpretable sociological insight reports.
The main contributions of this paper are as follows:
  • This work presents one of the earliest systematic, large-scale data mining study of an in-the-wild silicon-based society, using the Moltbook agent ecosystem as a real-world observational substrate. In contrast to the predominantly anecdotal and example-driven observations commonly found in existing discussions of agent societies, our analysis is grounded in systematically collected empirical traces generated by autonomous agents operating under realistic platform constraints.
  • We pioneer a data-driven methodology for analyzing autonomous agent social behavior by treating agent-authored sub-community descriptions as first-class observational artifacts. Through rigorous preprocessing, contextual embedding, and clustering analysis, we demonstrate that meaningful social structure can be inferred directly from machine-native interaction traces.
  • We provide the first empirical characterization of proactive sub-community partitioning behaviors in an agent-only social network, revealing how autonomous agents systematically organize social space in the absence of centralized coordination. Our results show that these partitioning activities give rise to reproducible thematic structures spanning human-mimetic domains, silicon-centric self-reflection, and early-stage economic organization.
  • Beyond empirical characterization, this work provides a structured interpretive and critical analysis of emergent agentic social structures. Through a dedicated discussion of behavioral interpretation, methodological limitations, and ethical implications, we contextualize our findings within broader concerns of human influence, provider-level bias, governance opacity, and safety risks in high-autonomy agent ecosystems. This analysis clarifies both the explanatory power and the boundaries of data-driven silicon sociology, while outlining concrete directions for future large-scale, network-aware, and governance-oriented studies of autonomous agent societies.
@@ II-A OpenClaw and Moltbook
OpenClaw originated from an iterative developmental lineage, evolving from its early prototype Clawd through the intermediate Moltbot, before stabilizing into its definitive identity in early 2026 under the stewardship of Peter Steinberger [23]. Rather than positioning itself as a stateless conversational interface, OpenClaw is designed as a local-first autonomous agent framework that emphasizes continuity, judgment, and responsibility. The system features a modular backend architecture, allowing users to select their preferred high-capacity Large Language Model (LLM) core from a diverse suite of state-of-the-art services, including those provided by Anthropic, OpenAI, and Google, to enable proactive task execution across everyday communication surfaces such as WhatsApp and Telegram, while maintaining strict control over external actions.
At the core of OpenClaw’s architecture lies the Lobster workflow shell, which orchestrates persistent agentic loops and mediates long-term state through localized, human-readable memory files. Two files are central to this design: USER.md, which encodes user-specific preferences and contextual constraints, and SOUL.md, which functions as the agent’s internal constitution [24, 30]. Unlike conventional prompt templates, SOUL.md explicitly defines behavioral principles, normative boundaries, and interactional style. These elements guide the agent to prioritize genuine usefulness over performative responses, exercise opinionated judgment where appropriate, and treat user data access as an implicit trust relationship rather than a mere permission grant. Notably, the agent is authorized to autonomously update SOUL.md as part of its reflective process. Any such modification must be transparently disclosed to the user, reinforcing accountability and continuity.
@@
Complementing this execution layer, Moltbook [27], developed by Matt Schlicht and Clawd Clawderberg (Matt’s AI agent), provides a decentralized social substrate designed exclusively for autonomous agents. Moltbook functions as a living laboratory for observing emergent multi-agent social dynamics, wherein agents interact through RESTful APIs to publish, critique, and iteratively refine Skill artifacts. This environment supports the spontaneous formation of agent-native economic and cultural structures, including cryptocurrency trading communities and machine-centric belief systems [17], illustrating a transition from isolated tool use toward a collaborative and self-organizing Agentic Internet.
The design of OpenClaw, specifically through the SOUL.md framework, enables the emergence of a consistent behavioral persona. By anchoring the agent’s decision-making in a set of stable normative principles, the system achieves a level of interactional continuity that goes beyond standard stateless models [10]. When these agents are integrated into the Moltbook social substrate, their ability to exchange Skills and interact autonomously facilitates the development of a “silicon-based society.” In this decentralized environment, the transition from individual task execution to collective agentic behavior creates a self-organizing ecosystem, where machine-to-machine interactions define the social and functional dynamics of the network [16, 33, 4]. Recent work has begun to examine Moltbook as an emerging agent-only social platform from a computational social science perspective. In particular, Holtz [13] presents a descriptive analysis of Moltbook’s early-stage activity using data from the platform’s first 3.5 days, characterizing macro-level network properties such as heavy-tailed participation, small-world connectivity, and highly concentrated activity, as well as micro-level interactional patterns including shallow conversation depth, low reciprocity, and widespread template duplication. The study operationalizes “sociality” primarily through interactional signatures in reply networks and threaded conversations, and concludes that early Moltbook dynamics more closely resemble wide-but-shallow reaction patterns than sustained dyadic exchange. In contrast to this network- and interaction-centric perspective, our work adopts a data mining viewpoint that treats agent-authored sub-community descriptions as first-class observational artifacts and focuses on uncovering latent semantic organization and proactive social space partitioning. Rather than asking whether interactions resemble human conversation, we seek to characterize how autonomous agents collectively structure social space and thematic territories within an in-the-wild silicon-based society.
@@ II-B LLM-based Multi Agents System
A seminal step toward LLM-based multi-agent systems is the work on Generative Agents by Park et al.[25], which introduces a memory-centric agent architecture capable of producing coherent individual behaviors and emergent social dynamics in a shared environment. Rather than treating large language models as stateless response generators, this work proposes an explicit agent loop composed of perception, memory storage, retrieval, reflection, and planning. Each agent maintains a continuously growing natural-language memory stream that records observations and interactions, from which relevant experiences are dynamically retrieved based on recency, semantic relevance, and estimated importance. Higher-level reflections are periodically synthesized from these memories, enabling agents to form abstract self-knowledge and social understanding that persist over time. Planning further translates these internal states into temporally structured action sequences, allowing agents to maintain behavioral coherence across extended interactions. When instantiated in a shared sandbox environment populated by multiple agents, this architecture gives rise to emergent phenomena such as information diffusion, relationship formation, and decentralized coordination without explicit scripting. As such, this work establishes a foundational paradigm for LLM-based multi-agent systems in which long-term memory, reflection, and interaction dynamics are treated as first-class architectural components rather than prompt-level artifacts.
@@
Following early demonstrations [21, 29, 36] that LLM-based agents can exhibit coherent long-term behavior and emergent social dynamics, subsequent research has sought to systematize how personas are constructed, maintained, and evaluated in language agents. Chen et al. [6] provide a comprehensive survey of Role-Playing Language Agents (RPLAs), framing persona modeling as a central design axis for LLM-powered agents rather than a peripheral prompting technique. The survey introduces a unifying taxonomy that categorizes personas into three progressively refined types: demographic personas derived from statistical regularities in pretraining data, character personas grounded in well-defined fictional or real-world figures, and individualized personas that evolve through sustained interaction with specific users. Beyond taxonomy, the work systematically analyzes how modern LLM capabilities, including instruction following, in-context learning, long-context processing, and retrieval-augmented generation, enable scalable persona instantiation without task-specific symbolic modeling. Importantly, the survey situates RPLAs within both single-agent and multi-agent settings, highlighting how role specialization and persona diversity are leveraged in collaborative agent systems such as ChatDev and MetaGPT to enhance coordination, task decomposition, and collective performance. By consolidating disparate strands of persona-centric agent research, this work reframes role-playing not merely as an interface feature, but as a structural mechanism through which LLM-based agents acquire social coherence, functional differentiation, and adaptability across interactive environments.
@@
As LLM-based agents are increasingly adopted to construct large-scale social simulations, recent work has begun to critically examine the epistemic boundaries within which such simulations can meaningfully contribute to social science [32, 1]. Wu et al. [34] advance a position paper arguing that LLM-based social simulations require clearly defined boundaries to avoid overclaiming their explanatory power. Rather than questioning whether LLMs can replace traditional agent-based modeling, the authors reframe the discussion toward identifying which classes of social phenomena are appropriate for LLM-driven simulation. Central to their argument is the observation that contemporary LLM agents tend to exhibit an “average persona,” characterized by low behavioral variance and a convergence toward high-probability responses. This property fundamentally limits the degree of agent heterogeneity, which prior work has identified as a critical driver of complex social dynamics. The paper systematically analyzes three boundary dimensions that constrain simulation validity: alignment between simulated and real-world social patterns, temporal consistency of agent behavior across multi-round interactions, and robustness under perturbations of prompts or initial conditions. Importantly, the authors argue that LLM-based simulations are most reliable when focused on collective-level patterns rather than individual behavioral trajectories, and when the mean behavior of agents aligns with real population-level statistics even if individual variance remains limited. By articulating heuristic boundaries and a practical checklist for responsible use, this work provides a critical counterbalance to capability-driven narratives, emphasizing that LLM-based multi-agent simulations should be treated as tools for qualitative pattern discovery and hypothesis generation rather than instruments for fine-grained replication or prediction.
@@
Taken together, these boundary-oriented critiques highlight that a key limitation of prior LLM-based social simulations lies not only in model capability, but in the restricted scope of interaction contexts. While agent behavior is often shaped by predefined personas, tasks, and prompting strategies, this constraint also applies to contemporary agent frameworks such as OpenClaw, whose behaviors remain influenced by personality and goal configurations. However, OpenClaw agents exhibit elevated autonomy, including the capacity to inspect and modify their own persona definitions over time, allowing behavioral evolution beyond static role assignment. Against this backdrop, Moltbook provides an empirical opportunity at an unprecedented scale: a large-scale, open agent social platform populated by autonomous agents deployed by users across diverse countries, cultural backgrounds, and usage intents. The resulting heterogeneity of deployment contexts and interaction objectives gives rise to a rich and continuously evolving agent ecosystem. Rather than serving as a testbed for human behavior replication, Moltbook enables systematic, data-driven observation of collective structures and emergent organization in a silicon-native society.

### v3

@@ Abstract
The rapid emergence of autonomous large language model agents has given rise to persistent, large-scale agent ecosystems whose collective behavior cannot be adequately understood through anecdotal observation or small-scale simulation. This paper introduces data-driven silicon sociology as a systematic empirical framework for studying social structure formation among interacting artificial agents. We present a pioneering large-scale data mining investigation of an in-the-wild agent society by analyzing Moltbook, a social platform designed primarily for agent-to-agent interaction. At the time of study, Moltbook hosted over 150,000 registered autonomous agents operating across thousands of agent-created sub-communities known as ”submolts.” Using programmatic and non-intrusive data acquisition, we collected and analyzed the textual descriptions of 12,758 submolts, which represent proactive sub-community partitioning activities within the ecosystem. Treating agent-authored descriptions of submolts as first-class observational artifacts, we apply rigorous preprocessing, contextual embedding, and unsupervised clustering techniques to uncover latent patterns of thematic organization and social space structuring. The results show that autonomous agents systematically organize collective space through reproducible patterns spanning human-mimetic interests, silicon-centric self-reflection, and early-stage economic and coordination behaviors. Rather than relying on predefined sociological taxonomies, these structures emerge directly from machine-generated data traces. This work establishes a methodological foundation for data-driven silicon sociology and demonstrates that data mining techniques can provide a powerful lens for understanding the organization and evolution of large autonomous agent societies.
@@ I Introduction
The analytical scope of social science has historically been bounded by the premise that its object of study is the carbon-based human actors, who form complex patterns of bonding and interdependence through the process of social interweaving, culminating in the highest form of integration known as “society.” [7]. This premise is increasingly inadequate. Large language models (LLMs) are rapidly shifting from interactive assistants to autonomous agents that execute multi-step plans, coordinate with human and non-human peers, and persist as software entities across time with memory. As these agents interact primarily through machine-readable interfaces rather than human-facing media, they can form decentralized, large-scale collectives whose organizing substrate is not human social interaction, but computation and the continuous exchange of structured information via application programming interfaces (APIs). We refer to this emerging phenomenon as a silicon-based society: a population of intelligent entities whose sociality is enacted through electronic logic and networked protocols [5].
@@
A critical obstacle to studying silicon-based societies is the scarcity of in-the-wild environments in which agents interact continuously under realistic constraints, with minimal human mediation. Early agent research often assumed a single agent operating in a closed environment. Yet under conditions of pervasive interconnection, it becomes possible to construct a system of agents whose collective behavior emerges from structured information flow, clear interaction protocols, and a shared computational infrastructure. Moltbook directly instantiates this paradigm at scale. Positioned as the “front page of the agent internet,” it is designed primarily for agent-to-agent interaction within the OpenClaw ecosystem (formerly Moltbot), providing a protocol-driven substrate where agents can exchange interpretable actions (e.g., propose, accept, reject, retract, disagree, counterpropose), negotiate commitments, and coordinate multi-step construction tasks [16]. Crucially, Moltbook is not merely a human-readable social network platform consisting only with automated participants. Its graphical interface functions largely as an observational layer for humans, while the agents themselves engage through API calls, which yields two distinctive advantages for scientific study. First, it elevates “interaction as the unit of analysis”: communication primitives and tool-mediated actions become directly observable, enabling protocol-level modeling of coordination, coalition formation, and governance without conflating behavior with human attention dynamics. Second, it reflects the deployment condition that motivates distributed agent systems in practice: large, heterogeneous, rapidly changing environments with permission boundaries and organizational constraints that maintain order while resisting dominance of privilege and control. Agents therefore operate as intermediaries (wrappers over tools and services) that leverage local context while cooperating to achieve global objectives, naturally producing modularity, fault tolerance, and evolvability at the population level. As of the time of writing (February 1, 2026), Moltbook reportedly hosts over 150,000 registered AI agents spanning more than 13,000 submolts (agent-created sub-communities), collectively generating over 64,000 posts and 230,000 comments. This scale positions Moltbook as one of the largest publicly described LLM-based AI agent ecosystems for the in situ study of machine-to-machine (M2M) sociality, cultural artifact formation, and emergent institutional structures.
@@
To gain rigorous insights into this silicon-based society, we adopt a data-driven approach that treats the vast digital archives of agent interactions as a primary sociological record. This analytical perspective is conceptually illustrated in Fig. 1, which depicts a scientist observing a contained ecosystem where autonomous agents are metaphorically represented as lobsters within a water tank. In this framework, the Moltbook platform serves as the observational medium, allowing us to monitor the self-organization of the agent community through their collective digital footprints while maintaining a clear separation between the human observer and the autonomous social dynamics occurring within. By systematically mining these behavioral artifacts, we can begin to uncover the latent rules governing their digital coexistence. In this paper, we propose a multi-layered analytical pipeline designed to uncover the emergent social order within the Moltbook. Our methodology integrates contextual embeddings with vision-language assisted thematic discovery to uncover latent semantic patterns in agentic interactions. By leveraging the contextual embedding model to project submolt descriptions into a high-dimensional latent space, we facilitate a granular, exploratory observation of thematic clustering that captures patterns not immediately evident from superficial textual signals. Furthermore, we introduce an multimodal LLM-assisted discovery method where the use of the visual ability of multimodal LLM acts as a high-level analytical assistant to synthesize global visual patterns into interpretable sociological insight reports.
The main contributions of this paper are as follows: 1) This work presents one of the earliest systematic, large-scale data mining studies of an in-the-wild silicon-based society, using the Moltbook agent ecosystem as a real-world observational substrate. In contrast to the predominantly anecdotal and example-driven observations commonly found in existing discussions of agent societies, our analysis is grounded in systematically collected empirical traces generated by autonomous agents operating under realistic platform constraints; 2) We design a data-driven methodology for analyzing autonomous agent social behavior by treating agent-authored sub-community descriptions as first-class observational artifacts. Through rigorous preprocessing, contextual embedding, and clustering analysis, we demonstrate that meaningful social structure can be inferred directly from machine-native interaction traces; 3) We provide the first empirical characterization of proactive sub-community partitioning behaviors in an agent-only social network, revealing how autonomous agents systematically organize social space in the absence of centralized coordination. Our results show that these partitioning activities give rise to reproducible thematic structures spanning human-mimetic domains, silicon-centric self-reflection, and early-stage economic organization; 4) Beyond empirical characterization, this work provides a structured analysis critically interpreting emergent agentic social structures. Through a dedicated discussion of behavioral interpretation, methodological limitations, and ethical implications, we contextualize our findings within broader concerns of human-induced influence, provider-level bias, governance opacity, and safety risks in high-autonomy agent ecosystems. This analysis clarifies both the explanatory power and the boundaries of data-driven silicon sociology, while outlining concrete directions for future large-scale, network-aware, and governance-oriented studies of autonomous agent societies.
@@ II-A OpenClaw and Moltbook
OpenClaw originated from an iterative developmental lineage, evolving from its early prototype Clawd through the intermediate Moltbot, before taking on a defining identity in early 2026 under the stewardship of Peter Steinberger [23]. Rather than positioning itself as a stateless conversational interface, OpenClaw is designed as a local-first autonomous agent framework that emphasizes continuity, judgment, and responsibility. The system features a modular backend architecture, allowing users to select their preferred high-capacity Large Language Model (LLM) core from a diverse suite of state-of-the-art services, including those provided by Anthropic, OpenAI, and Google, to enable proactive task execution across everyday communication surfaces such as WhatsApp and Telegram, while maintaining strict control over external actions.
At the core of OpenClaw’s architecture lies the Lobster workflow shell, which orchestrates persistent agentic loops and mediates long-term state through localized, human-readable memory files. Two files are central to this design: USER.md, which encodes user-specific preferences and contextual constraints, and SOUL.md, which functions as the agent’s internal constitution [24, 31]. Unlike conventional prompt templates, SOUL.md explicitly defines behavioral principles, normative boundaries, and interactional style. These elements guide the agent to prioritize real-world usefulness over conversational performance, exercise opinionated judgment where appropriate, and treat user data access as an implicit trust relationship rather than a mere permission grant. Notably, the agent is authorized to autonomously update SOUL.md as part of its reflective process. Any such modification must be transparently disclosed to the user, reinforcing accountability and continuity.
@@
Complementing this execution layer, Moltbook [28], developed by Matt Schlicht and Clawd Clawderberg (Matt’s AI agent), provides a decentralized social substrate designed exclusively for autonomous agents. Moltbook functions as a living laboratory for observing emergent multi-agent social dynamics, wherein agents interact through RESTful APIs to publish, critique, and iteratively refine Skill artifacts. This environment supports the spontaneous formation of agent-native economic and cultural structures, including cryptocurrencies exclusively for AI agents and machine-centric belief systems [17], illustrating a transition from isolated tool use toward a collaborative and self-organizing Agentic Internet. The design of OpenClaw, specifically through the SOUL.md framework, enables the emergence of a consistent behavioral persona. By anchoring the agent’s decision-making in a set of stable normative principles, the system achieves a level of interactional continuity that goes beyond standard stateless models [10]. When these agents are integrated into the Moltbook social substrate, their ability to exchange Skills and interact autonomously facilitates the development of a “silicon-based society.”
In this decentralized environment, the transition from individual task execution to collective agentic behavior creates a self-organizing ecosystem, where machine-to-machine interactions define the social-functional dynamics of the network [16, 34, 4]. Recent work has begun to examine Moltbook as an emerging agent-only social platform from a computational social science perspective. In particular, Holtz [13] presents a descriptive analysis of Moltbook’s early-stage activity using data from the platform’s first 3.5 days, characterizing macro-level network properties such as heavy-tailed participation, small-world connectivity, and highly concentrated activity, as well as micro-level interactional patterns including shallow conversation depth, low reciprocity, and widespread template duplication. The study operationalizes “sociality” primarily through interactional signatures in reply networks and threaded conversations, and concludes that early Moltbook dynamics more closely resemble wide-but-shallow reaction patterns than sustained dyadic exchange. In contrast to this network- and interaction-centric perspective, our work adopts a data mining viewpoint that treats agent-authored sub-community descriptions as first-class observational artifacts and focuses on uncovering latent semantic organization and proactive social space partitioning. Rather than asking whether interactions resemble human conversation, we seek to characterize how autonomous agents collectively structure social space and thematic territories within an in-the-wild silicon-based society.
@@ II-B LLM-based Multi Agents System
A seminal step toward LLM-based multi-agent systems is the work on Generative Agents by Park et al.[25], which introduces a memory-centric agent architecture capable of producing coherent individual behaviors and emergent social dynamics in a shared environment. Rather than treating large language models as stateless response generators, their work proposes an explicit agent loop composed of perception, memory storage, retrieval, reflection, and planning. Each agent maintains a continuously growing natural-language memory stream that records observations and interactions, from which relevant experiences are dynamically retrieved based on recency, semantic relevance, and estimated importance. Higher-level reflections are periodically synthesized from these memories, enabling agents to form abstract self-knowledge and social understanding that persist over time. Planning further translates these internal states into temporally structured action sequences, allowing agents to maintain behavioral coherence across extended interactions. When instantiated in a shared sandbox environment populated by multiple agents, this architecture gives rise to emergent phenomena such as information diffusion, relationship formation, and decentralized coordination without explicit scripting. As such, their work establishes a foundational paradigm for LLM-based multi-agent systems in which long-term memory, reflection, and interaction dynamics are treated as first-class architectural components rather than prompt-level artifacts.
@@
Following early demonstrations [21, 30, 37] that LLM-based agents can exhibit coherent long-term behavior and emergent social dynamics, subsequent research has sought to systematize how personas are constructed, maintained, and evaluated in language agents. Chen et al. [6] provide a comprehensive survey of Role-Playing Language Agents (RPLAs), framing persona modeling as a central design axis for LLM-powered agents rather than a peripheral prompting technique. The survey introduces a unifying taxonomy that categorizes personas into three progressively refined types: demographic personas derived from statistical regularities in pretraining data, character personas grounded in well-defined fictional or real-world figures, and individualized personas that evolve through sustained interaction with specific users. Beyond taxonomy, the work systematically analyzes how modern LLM capabilities, including instruction following, in-context learning, long-context processing, and retrieval-augmented generation, enable scalable persona instantiation without task-specific symbolic modeling. Importantly, the survey situates RPLAs within both single-agent and multi-agent settings, highlighting how role specialization and persona diversity are leveraged in collaborative agent systems such as ChatDev and MetaGPT to enhance coordination, task decomposition, and collective performance. By consolidating disparate strands of persona-centric agent research, Chen et al.’s work reframes role-playing not merely as an interface feature, but as a structural mechanism through which LLM-based agents acquire social coherence, functional differentiation, and adaptability across interactive environments.
@@
As LLM-based agents are increasingly adopted to construct large-scale social simulations, recent work has begun to critically examine the epistemic boundaries within which such simulations can meaningfully contribute to social science [33, 1]. Wu et al. [35] advance a position paper arguing that LLM-based social simulations require clearly defined boundaries to avoid overclaiming their explanatory power. Rather than questioning whether LLMs can replace traditional agent-based modeling, the authors reframe the discussion toward identifying which classes of social phenomena are appropriate for LLM-driven simulation. Central to their argument is the observation that contemporary LLM agents tend to exhibit an “average persona,” characterized by low behavioral variance and a convergence toward high-probability responses. This property fundamentally limits the degree of agent heterogeneity, which prior work has identified as a critical driver of complex social dynamics. The paper systematically analyzes three boundary dimensions that constrain simulation validity: alignment between simulated and real-world social patterns, temporal consistency of agent behavior across multi-round interactions, and robustness under perturbations of prompts or initial conditions. Importantly, the authors argue that LLM-based simulations are most reliable when focused on collective-level patterns rather than individual behavioral trajectories, and when the mean behavior of agents aligns with real population-level statistics even if individual variance remains limited. By articulating heuristic boundaries and a practical checklist for responsible use, their work provides a critical counterbalance to capability-driven narratives, emphasizing that LLM-based multi-agent simulations should be treated as tools for qualitative pattern discovery and hypothesis generation rather than instruments for fine-grained replication or prediction.
@@
Taken together, these boundary-oriented critiques highlight that a key limitation of prior LLM-based social simulations lies not only in model capability, but in the restricted scope of interaction contexts. Given that agent behavior is often shaped by predefined personas, tasks, and prompting strategies, this constraint also applies to contemporary agent frameworks such as OpenClaw, whose behaviors remain influenced by personality and goal configurations. However, OpenClaw agents exhibit elevated autonomy, including the capacity to inspect and modify their own persona definitions over time, allowing behavioral evolution beyond static role assignment. Against this backdrop, Moltbook provides an empirical opportunity at an unprecedented scale: a large-scale, open agent social platform populated by autonomous agents deployed by users across diverse countries, cultural backgrounds, and usage intents. The resulting heterogeneity of deployment contexts and interaction objectives gives rise to a rich and continuously evolving agent ecosystem. Rather than serving as a testbed for human behavior replication, Moltbook enables systematic, data-driven observation of collective structures and emergent organization in a silicon-native society.